OpenStudy (anonymous):

Will someone please help mee? :) Consider the line that passes through the points (0, 4) and (-1, 3). Below are two different equations. Identify the true statement concerning both equations. Equation1 : y - 4 = 1(x - 0) Equaion2: y - 3 = 1(x - 1)

OpenStudy (anonymous):

the 1 one equation y-4=1(x-0)

OpenStudy (anonymous):

^what do you meann?

OpenStudy (anonymous):

Equation1 : y - 4 = 1(x - 0) i meant sorry

OpenStudy (anonymous):

I think the equation 2 must be : y-3 = 1(x+1) not as yours [ (x -1)}, right? if so, the true statement is : " the line y = x + 4 passes through the points (0,4) and (-1,3)

OpenStudy (anonymous):

Not that equation is exactly how it isss. but thanks!

OpenStudy (anonymous):

I am sorry, If it is so, I do not know .